# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

### v4

@@ 7 Conclusion
As disruptions are ubiquitous in urban transportation systems and are expected to increase continuously with population growth and ongoing urbanization, this work introduces the novel concept of antifragility into traffic control to tackle the growing trend of disruptions and the fragile nature of road transportation networks. First, from a terminology perspective, antifragility is compared with other related concepts commonly applied in transportation, including robustness, resilience, reliability, and adaptiveness. Through reviewing previous research on applying RL algorithms in traffic control to achieve robust or resilient traffic control systems, we explore whether and how these properties can be induced with RL. Building on such concepts, we aim to induce antifragility in perimeter control by modifying the state definition and reward function based on a state-of-the-art RL-based perimeter control algorithm. We incorporate the first and second derivatives, representing the change rate and the curvature of traffic states, to provide the RL algorithm with richer and reliable information under disruptive events. To mitigate the potential oscillations caused by integrating derivatives, a damping term is introduced to stabilize the computed actions. Furthermore, a redundant overcompensation term in the reward function strengthens the system’s antifragility against disruptions. In addition to the performance gain and antifragility, another contribution of this work is the consideration of real-world observability constraints, complementing the drawback of the current state-of-the-art algorithms of utilizing hardly observable measurements.
@@
We conducted comprehensive experiments to compare our proposed antifragile RL-based perimeter control approach against three other methods: MPC, MPC-MHE, and an RL-based algorithm. Two distinct manifestations of disruptions are examined, i.e., demand and supply disruptions. Uncertainties in the disruption magnitude are also considered. The results first demonstrate the effectiveness of our proposed antifragile RL-based algorithm, through its superior performance in terms of lower TTS. More importantly, by putting forward a novel method for quantifying antifragility through computing its performance distribution skewness, we confirm that the proposed algorithm exhibits greater antifragility with the lowest skewness among all the methods examined, delivering increasingly better performance as the magnitude of disruptions grows. Finally, the performance and antifragile properties are studied with limited real-world observability, showing that a certain degree of performance can be traded for using only accessible sensor measurements in reality.
@@
Several limitations of the current study need to be addressed as well. For instance, measurements and model uncertainties are commonly studied when designing perimeter control algorithms to validate their robustness. However, to uphold the clarity of this work, that is, studying antifragility instead of robustness, and also due to our introduction of uncertainties across the episodes with incremental disruptions, we disregard the uncertainties within each episode, but we acknowledge the common presence of uncertainties in the real urban networks. Furthermore, as this work primarily works on numerical simulation, the real-world operation together with the proposed antifragile RL-based algorithm can be better validated with microsimulation software, such as SUMO, which is subject to our future work.
@@
In conclusion, this study is the first of its kind to pioneer the application of antifragility in the operation of transportation systems, continuously improving the performance of a network under unforeseen disruptions with learning-based algorithms. Additionally, it introduces a new paradigm for evaluating system operation under disruptive conditions. Moreover, the concept is generic to be extended to other traffic control systems and potentially to any system subject to growing disruptions.
@@ 8 CRediT authorship contribution statement
Linghang Sun: Conceptualization, Investigation, Methodology, Visualization, Writing – original draft. Michail A. Makridis: Conceptualization, Methodology, Supervision, Writing - review & editing. Alexander Genser: Methodology, Visualization, Writing - review & editing. Cristian Axenie: Project administration, Resources, Writing - review & editing. Margherita Grossi: Project administration, Resources, Writing - review & editing. Anastasios Kouvelas: Supervision, Writing - review & editing.
@@ 9 Acknowledgement
This research was kindly funded by the Huawei Munich Research Center under the framework of the Antigones project.
@@ Appendix A Graphical illustration of concepts related to antifragility
Figure 16: Illustration for robustness, resilience, adaptiveness, and antifragility

### v5

@@ 7 Conclusion
As disruptions are ubiquitous in urban transportation systems and are expected to increase continuously with population growth and ongoing urbanization, this work introduces the novel concept of antifragility into traffic control to tackle the growing trend of disruptions and the fragile nature of road transportation networks. First, from a terminology perspective, antifragility is compared with other related concepts commonly applied in transportation, including robustness, resilience, reliability, and adaptiveness. Through reviewing previous research on applying RL algorithms in traffic control to achieve robust, resilient, or reliable systems from either the reliability engineering or traffic control perspective, we explore whether and how these properties can be induced with RL. Building on such concepts, we aim to induce antifragility in perimeter control by modifying the state definition and reward function based on a state-of-the-art RL-based perimeter control algorithm. We incorporate the first and second derivatives, representing the change rate and the curvature of traffic states, to provide the RL algorithm with richer and reliable information under disruptive events. To mitigate the potential oscillations caused by integrating derivatives, a damping term is introduced to stabilize the computed actions. Furthermore, a redundant overcompensation term in the reward function strengthens the system’s antifragility against disruptions. In addition to the performance gain and antifragility, another contribution of this work is the consideration of real-world observability constraints, complementing the drawback of the current algorithms of utilizing hardly observable measurements.
@@
We conducted comprehensive experiments to compare the proposed antifragile RL-based perimeter control approach against three other methods: MPC, MPC-MHE, and an RL-based algorithm. Two distinct manifestations of disruptions are examined, i.e., demand and supply disruptions. Uncertainties in the disruption magnitude are also considered. The results first demonstrate the effectiveness of our proposed antifragile RL-based algorithm, through its superior performance in terms of lower TTS. More importantly, by putting forward a novel method for quantifying antifragility through computing its performance distribution skewness, we confirm that the proposed algorithm exhibits greater antifragility with the lowest skewness among all the methods examined, delivering increasingly better performance as the magnitude of disruptions grows. Moreover, the performance and antifragile properties are studied with real-world limited observability, showing that a certain degree of performance can be traded for using only accessible sensor measurements in reality. We further validated the superior performance and relative antifragile properties of the proposed algorithm with a real-world case study. Realistic demand, supply, and disruptions are created from multiple sources of data gathered in the city center of Zurich. At last, since the proposed methodology can function as an add-on to other contemporary RL algorithms, several more advanced baselines, i.e., TD3 and SAC, are also involved. Results demonstrated the broad applicability of our proposed approach to other contemporary baselines.
@@
Several limitations of the current study need to be addressed as well. For instance, this work primarily studies numerical simulation, with Section 6.3 showcasing a numerical simulation created from real-world network, demand, and disruption data. It should be acknowledged that real-world operation with the proposed antifragile RL-based algorithm can be more accurately validated through traffic microsimulation software like SUMO. However, due to the complexity of selecting one or multiple functional strings of traffic lights to form the perimeter in a complicated real-world network, such works are generally carried out as individual research, as in Kampitakis and Vlahogianni (2025), which is listed as one of the future directions for the present study. Furthermore, measurements and model uncertainties are commonly studied when designing perimeter control algorithms to validate their robustness. However, to uphold the clarity of this work, that is, studying antifragility instead of robustness, and also due to our introduction of uncertainties across the episodes with incremental disruptions, we only studied such time-varying demand uncertainties in the case study created from real-world datasets.
@@
In conclusion, this study is the first of its kind to pioneer the application of antifragility in the operation and control of transportation systems, continuously improving the performance of a network under unforeseen disruptions with learning-based algorithms. Additionally, it introduces a new paradigm for evaluating system operation under disruptive conditions through computing the distribution skewness. Moreover, the concept is well generic to be extended to other traffic control systems and potentially to control systems in general that are subject to growing disruptions.
@@ 8 CRediT authorship contribution statement
Linghang Sun: Conceptualization, Investigation, Methodology, Visualization, Writing – original draft. Michail A. Makridis: Conceptualization, Methodology, Supervision, Writing - review & editing. Alexander Genser: Methodology, Visualization, Writing - review & editing. Cristian Axenie: Conceptualization, Project administration, Resources, Writing - review & editing. Margherita Grossi: Project administration, Resources, Writing - review & editing. Anastasios Kouvelas: Supervision, Writing - review & editing.
@@ 9 Acknowledgement
This research was primarily funded by the Huawei Munich Research Center through the Antigones project. Additional financial support has been received from the Swiss State Secretariat for Education, Research, and Innovation (SERI) with contract number 25.00120, as part of the European Union’s Horizon Europe project AntifragiCity with project ID 101203052.
@@ Appendix A Graphical illustration of concepts related to antifragility
Figure 12: Illustration for robustness, resilience, adaptiveness, and antifragility.